# NaturalSpeech 3: Zero-Shot Speech Synthesis with Factorized Codec and Diffusion Models

### v1

@@ B.2 Reconstruction Performance Comparison
We evaluate the reconstruction performance with the following objective metrics: Perceptual Evaluation of Speech Quality (PESQ), Short-Time Objective Intelligibility (STOI), Multi-Resolution STFT Distance (MSTFT), and Mel-Cepstral Distortion (MCD). These metrics collectively measure the difference between the original and the reconstructed samples. We select the following open-source codec models as baselines: EnCodec [15] [^8] https://github.com/facebookresearch/encodec, HiFi-Codec [63] [^9] https://github.com/yangdongchao/AcademiCodec, and Descript-Audio-Codec (DAC) [32] [^10] https://github.com/descriptinc/descript-audio-codec. We additionally reproduce SoundStream [14] following the original paper’s implementation and experimental setup. Table  12 shows that 1) FACodec significantly surpasses SoundStream in the same bandwidth setting ($0.44$ in PESQ, $0.05$ in STOI, $0.14$ in MSTFT and $0.79$ in MCD, respectively). Moreover, FACodec achieves on-par performance with SoundStream even when its bandwidth is doubled ($0.02$ in PESQ, $0.01$ in STOI, $-0.01$ in MSTFT and $0.17$ in MCD, respectively). 2) For a fair comparison, we compare FACodec with other baselines in a similar bandwidth. FACodec achieve comparable or better result on most metrics than these strong baselines, which means that we can still achieve excellent reconstruction speech quality when disentangling speech attributes.
Table 12: The reconstruction quality evaluation of codecs. ${}^{\clubsuit}$ means results are infered from offical checkpoints. ${}^{\bigstar}$ means the reproduced checkpoint. means the reproduced model following the original paper’s implementation and experimental setup. All models use a codebook size of $1024$. We use bold to indicate the best result and underline to indicate the second-best result. Abbreviation: H (Hop Size), N (Codebook Number).
@@
${}^{\clubsuit}$ 24kHz 320 8 6.0 kbps 3.28 0.94 0.99 2.70
${}^{\bigstar}$ 16kHz 320 10 5.0 kbps 3.10 0.92 0.97 3.10
${}^{\clubsuit}$ 16kHz 320 4 2.0 kbps 3.17 0.93 0.98 3.05
${}^{\clubsuit}$ 16kHz 320 9 4.5 kbps 3.52 0.95 0.97 2.65

### v2

@@ B.2 Reconstruction Performance Comparison
We evaluate the reconstruction performance with the following objective metrics: Perceptual Evaluation of Speech Quality (PESQ), Short-Time Objective Intelligibility (STOI), Multi-Resolution STFT Distance (MSTFT), and Mel-Cepstral Distortion (MCD). These metrics collectively measure the difference between the original and the reconstructed samples. We select the following open-source codec models as baselines: EnCodec [15] [^11] https://github.com/facebookresearch/encodec, HiFi-Codec [65] [^12] https://github.com/yangdongchao/AcademiCodec, and Descript-Audio-Codec (DAC) [32] [^13] https://github.com/descriptinc/descript-audio-codec. We additionally reproduce SoundStream [14] following the original paper’s implementation and experimental setup. Table  13 shows that 1) FACodec significantly surpasses SoundStream in the same bandwidth setting ($0.44$ in PESQ, $0.05$ in STOI, $0.14$ in MSTFT and $0.79$ in MCD, respectively). Moreover, FACodec achieves on-par performance with SoundStream even when its bandwidth is doubled ($0.02$ in PESQ, $0.01$ in STOI, $-0.01$ in MSTFT and $0.17$ in MCD, respectively). 2) For a fair comparison, we compare FACodec with other baselines in a similar bandwidth. FACodec achieve comparable or better result on most metrics than these strong baselines, which means that we can still achieve excellent reconstruction speech quality when disentangling speech attributes.
Table 13: The reconstruction quality evaluation of codecs. means results are infered from offical checkpoints. means the reproduced checkpoint. means the reproduced model following the original paper’s implementation and experimental setup. All models use a codebook size of $1024$. We use bold to indicate the best result and underline to indicate the second-best result. Abbreviation: H (Hop Size), N (Codebook Number).
@@
EnCodec 24kHz 320 8 6.0 kbps 3.28 0.94 0.99 2.70
EnCodec 16kHz 320 10 5.0 kbps 3.10 0.92 0.97 3.10
HiFi-Codec 16kHz 320 4 2.0 kbps 3.17 0.93 0.98 3.05
DAC 16kHz 320 9 4.5 kbps 3.52 0.95 0.97 2.65